# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v2

@@ V Conclusion
In this letter, we have derived a strict isomorphism between a minimalist active matter model and resistively shunted junction theory, by recovering the Adler equation in the model’s dynamics (Eq. 9), and subsequently following the formalism developed by Ref. [11] for individual disordered Josephson junctions. We demonstrated that the system is governed by a tilted washboard potential, where a thermodynamic transition separates resistive ‘running’ agents from the ‘trapped’ population, the latter of which explicitly sustain an information supercurrent. As such, we have resolved the origin of emergent inertial flows in overdamped chiral active matter [2]. This establishes the minimalist active matter model as a distributed, disordered, motile Josephson array.

### v3

@@ VII Conclusion
We have derived a strict isomorphism between a minimalist, two-dimensional chiral active matter model, and resistively shunted junction theory, by recovering the Adler equation in the model’s dynamics (Eq. 9). We subsequently followed the formalism developed by Ref. [22] for individual disordered Josephson junctions, thereby demonstrating that the system is governed by a tilted washboard potential. A thermodynamic transition separates resistive ‘running’ agents from the ‘trapped’ population, the latter of which explicitly sustain an information supercurrent.
By extending the model to three dimensions, we recovered the Landau-Lifshitz-Gilbert equation for an overdamped spintronic fluid. The information supercurrents that we have predicted and described in empirical terms transmit phase gradients across the chiral flock, sustaining Goldstone modes from the precession around the mean field of the 3D active bath. The resulting spin waves advect phase gradients, providing inertia to the inviscid Euler turbulence of the Onsager dipole. The “phantom inertia” discovered in Ref. [8], $\lambda=R^{2}$, is therefore physically identified as the precession stiffness of the order parameter in three dimensions.
If substantiated in future forays, the transition to active chiral turbulence must be considered the emergence of a topological quasiparticle gas, where macroscopic Onsager dipoles, in turn bound states of phase singularities, acquire effective inertia from the Goldstone modes (azimuthal precession) of the underlying broken symmetry.
@@ Acknowledgements
Appendix A: Deriving the dispersion relation
We consider the spin waves supported by the azimuthal turning inertia and perturb the ordered state, linearizing the polarized state around small perturbations to $\hat{n}(\mathbf{r},t)$. That is, a flock perfectly aligned along the $z$-axis with a small perturbation $\mathbf{m}$ in the $xy$-plane:
$\hat{n}(\mathbf{r},t)\approx\hat{z}+\mathbf{m}(\mathbf{r},t)\quad\text{where}\quad\mathbf{m}=(m_{x},m_{y},0),$ (34)
which we shall eventually insert into Eq. (30).
Neglecting higher-order terms ($|\mathbf{m}|^{2}\approx 0$), the Laplacian is $\nabla^{2}\hat{n}\approx(\nabla^{2}m_{x},\nabla^{2}m_{y},0)$. We now evaluate the cross products to first order in $\mathbf{m}$:
$\hat{z}\times\nabla^{2}\mathbf{m}=\nabla^{2}m_{x}\hat{y}-\nabla^{2}m_{y}\hat{x}$, (35)
and,
$\hat{z}\times(\hat{z}\times\nabla^{2}\mathbf{m})=-\nabla^{2}\mathbf{m}.$ (36)
The former corresponds to inertial precession, while the latter leads to alignment. Substituting these expressions back into the extended equation:
$\frac{\partial\mathbf{m}}{\partial t}=-\frac{a_{0}\xi^{2}}{R}(\nabla^{2}m_{x}\hat{y}-\nabla^{2}m_{y}\hat{x})+(a_{0}R\xi^{2})\nabla^{2}\mathbf{m}.$ (37)
Separating into vector components yields:
$\dot{m}_{x}$ $=\frac{a_{0}\xi^{2}}{R}\nabla^{2}m_{y}+(a_{0}R\xi^{2})\nabla^{2}m_{x},$ (38)
$\dot{m}_{y}$ $=-\frac{a_{0}\xi^{2}}{R}\nabla^{2}m_{x}+(a_{0}R\xi^{2})\nabla^{2}m_{y}.$ (39)
Proposing a plane wave solution $\mathbf{m}\sim e^{i(kx-\omega t)}$ (with $\partial_{t}\to-i\omega$ and $\nabla^{2}\to-k^{2}$), the coupled equations read:
$-i\omega m_{x}$ $=-\frac{a_{0}\xi^{2}}{R}k^{2}m_{y}-(a_{0}R\xi^{2})k^{2}m_{x},$ (40)
$-i\omega m_{y}$ $=+\frac{a_{0}\xi^{2}}{R}k^{2}m_{x}-(a_{0}R\xi^{2})k^{2}m_{y}.$ (41)
Rearranging into matrix form $\mathbf{M}\mathbf{m}=0$:
$\begin{pmatrix}-i\omega+(a_{0}R\xi^{2})k^{2}&\frac{a_{0}\xi^{2}}{R}k^{2}\\ -\frac{a_{0}\xi^{2}}{R}k^{2}&-i\omega+(a_{0}R\xi^{2})k^{2}\end{pmatrix}\begin{pmatrix}m_{x}\\ m_{y}\end{pmatrix}=0.$ (42)
For non-trivial solutions, the determinant must be zero:
$(-i\omega+a_{0}R\xi^{2}k^{2})^{2}+\left(\frac{a_{0}\xi^{2}}{R}k^{2}\right)^{2}=0.$ (43)
Solving for $\omega$:
$i\omega+a_{0}R\xi^{2}k^{2}=\mp i\frac{a_{0}\xi^{2}}{R}k^{2},$ (44)
$\omega(k)=\pm\frac{a_{0}\xi^{2}}{R}k^{2}-i(a_{0}R\xi^{2})k^{2}.$ (45)
Absorbing the interaction scale $\xi^{2}$ into the normalization (or setting $\xi=1$), we recover the final dispersion relation:
$\omega(k)=\pm\frac{a_{0}}{R}k^{2}-i(a_{0}R)k^{2}.$ (46)